Wardell Company purchased a mainframe on January 1, 2019, at a cost of $48,000. The computer was depreciated using the straight-line method over an estimated five-year life with an estimated residual value of $3,000. On January 1, 2021, the estimate of useful life was changed to a total of 10 years, and the estimate of residual value was changed to $1,200.

2. Prepare the year-end journal entry for depreciation in 2021. Assume that the company uses the sum-of-the-years' -digits method instead of the straight-line method.
Answer #1

Sum of year digit = 5+4+3+2+1 = 15

Accumulated depreciation = (48000-3000)9/15 = 27000

2021 Dep = (48000-27000-1200)*8/36 = 4400

Account and explanation Debit Credit
Depreciation expense 4400
Accumulated depreciation-Computer 4400